DickT16:

ALGEBRA 2:Solve the following system of equations and show all work. y = −x2 + 4 y = 2x + 1

danielfootball123:

the ordered pairs are (-3, -5) and (1, 3).

bryanferia:

x=1 or x=-3 y=3 y=-5

danielfootball123:

you there

DickT16:

? how do you do it?

bryanferia:

@Crackert16 wrote:
? how do you do it?
what do you mean

bryanferia:

oh ok sorry

DickT16:

How do you solve it

bryanferia:

i will come back with the steps ok

DickT16:

Alright thank you

bryanferia:

do you mean x^2

danielfootball123:

thats impossible

danielfootball123:

to slove

danielfootball123:

trust me i know

DickT16:

I just mean how do you solve it step for step I want to learn and i need to solve it. And i do you know?

bryanferia:

@danielfootball123 wrote:
trust me i know
thank why is it a math question I don't think teachers would give an impossible question

danielfootball123:

y=-x2 + 4, y=2x+1 Solution: y = -x2 + 4 y = 2x + 1 These linear equations can be solved by using elimination or substitution method As both the equations are equal to y, the two equations are equal to one another. -x2 + 4 = 2x + 1 By further calculation 4 - 1 = 2x + x2 It can be written as x2 + 2x = 3 x2 + 2x - 3 = 0 (x + 3) (x - 1) = 0 So we get x = -3 or 1 Substituting the value of x in either of the equations y = 2x + 1 or y = 2x + 1 y = 2 (-3) + 1 or y = 2 (1) + 1 y = - 6 + 1 or y = 2 + 1 y = -5 or y = 3 Therefore, the ordered pairs are (-3, -5) and (1, 3). Solve the following system of equations and show all work. Summary: Solving the following system of equations, the ordered pairs are (-3, -5) and (1, 3).

danielfootball123:

sorry

DickT16:

me neither. So what are the steps bryan?

danielfootball123:

here you go

danielfootball123:

y=-x2 + 4, y=2x+1 Solution: y = -x2 + 4 y = 2x + 1 These linear equations can be solved by using elimination or substitution method As both the equations are equal to y, the two equations are equal to one another. -x2 + 4 = 2x + 1 By further calculation 4 - 1 = 2x + x2 It can be written as x2 + 2x = 3 x2 + 2x - 3 = 0 (x + 3) (x - 1) = 0 So we get x = -3 or 1 Substituting the value of x in either of the equations y = 2x + 1 or y = 2x + 1 y = 2 (-3) + 1 or y = 2 (1) + 1 y = - 6 + 1 or y = 2 + 1 y = -5 or y = 3 Therefore, the ordered pairs are (-3, -5) and (1, 3). Solve the following system of equations and show all work. Summary: Solving the following system of equations, the ordered pairs are (-3, -5) and (1, 3).